### v2

# Erasing Without Remembering: Implicit Knowledge Forgetting in Large Language Models

Abstract
In this paper, we investigate knowledge forgetting in large language models with a focus on its generalisation—ensuring that models forget not only specific training samples but also related implicit knowledge. To this end, we begin by identifying a broader unlearning scope that includes both target data and logically associated samples, including rephrased, subject-replaced, one-hop reasoned, and relation-reversed data. To rigorously evaluate generalisation, we introduce UGBench—the first comprehensive benchmark specifically designed to assess the unlearning of in-scope implicit knowledge covering 13 state-of-the-art methods across three datasets. UGBench reveals that unlearned models can still recall paraphrased answers and retain target facts in intermediate layers. This motivates us to take a preliminary step toward more generalised implicit knowledge forgetting by proposing PerMU—a novel probability perturbation-based unlearning paradigm. PerMU simulates adversarial unlearning samples to eliminate fact-related tokens from the logit distribution, collectively reducing the probabilities of all answer-associated tokens. Experiments are conducted on a diverse range of datasets, including TOFU, Harry Potter, ZsRE, WMDP, and MUSE, using models ranging from 1.3B to 13B in scale. The results demonstrate that PerMU delivers up to a 50.40% improvement in unlearning vanilla target data while maintaining a 40.73% boost in forgetting implicit knowledge. Our code can be found in https://github.com/MaybeLizzy/UGBench.
Index Terms:
Machine Unlearning, Large Language Models, Generalisation.
1 Introduction
Large language models (LLMs) [1, 2], while displaying remarkable performance thanks to their capacity for recalling extensive knowledge from pre-training corpora, are also increasingly susceptible to generating private, harmful, or even illegal content, due to their unintended memorisation of confidential information [3, 4]. In response to this dilemma, LLM-tailored machine unlearning [5, 6] has emerged as a rising research focus, aiming to develop reliable and computationally efficient knowledge-forgetting approaches for erasing the influence of specific undesired data from trained LLMs, all while preserving their utility for the remaining data.
Figure 1: Depiction of the proposed unlearning scope in a hypothetical semantic embedding space, highlighting the generalisation dilemma inherent in machine unlearning for LLMs. Ideally, hard in-scope samples that lie within the unlearning scope by a small margin should also be forgotten. These include rephrased questions, as well as the relation reversed questions and so on.
State-of-the-art machine unlearning approaches for LLMs broadly fall into two categories: training-free and training-based methods. The former, such as neuron editing [7], in-context learning [8], and prompt engineering [9], unlearn knowledge without additional training but often suffer from limited application scenarios. In contrast, training-based methods typically achieve greater unlearning effectiveness by updating model parameters and gradients, using techniques like gradient ascent [10], preference optimisation [11], relabeling-based fine-tuning [12], task arithmetic [13], logit-difference fine-tuning [14], or adding new parameters [15, 16].
Despite significant progress in LLM-based unlearning, this paper identifies an embarrassingly simple yet critical dilemma: existing methods typically teach models to forget only the exact expressions of the unlearning samples, while failing to genuinely unlearn paraphrased or other related information that should also be erased. To better illustrate, we introduce an unlearning scope in Fig. 1, encompassing all the knowledge that unlearned models are expected to forget, such as paraphrased versions, reversed relations, one-hop questions, and those with substituted subjects. As shown in Fig. 1, successful unlearning should intuitively modify the model’s behavior for in-scope samples while leaving out-of-scope samples unaffected [17]. However, to preserve utility, existing methods often compromise by achieving superficial forgetting, failing to unlearn “hard” samples near the boundary of the unlearning scope, which ultimately results in poor generalisation.
To further substantiate this observation, our first contribution in this paper is to introduce UGBench, the first comprehensive benchmark specifically designed to evaluate the unlearning ability of existing methods in forgetting neighbouring implicit knowledge. The benchmark spans three data domains: two widely-used machine unlearning datasets, TOFU [18] and the Harry Potter books [19, 20], as well as a popular model editing dataset, ZsRE [21]. Evaluations are performed on 13 existing methods using two language models of different scales, Phi-1.3B and LLaMA2-7B. Our empirical analyses on UGBench uncover the following unique findings that have been overlooked by existing research:
  • Identified challenge: Existing machine unlearning methods consistently exhibit a lack of generalisation;
[…]
These insights suggest that the target knowledge is not fully erased, leading to generalisation failure.
Motivated by the findings from UGBench, we highlight a critical challenge in LLM unlearning: enhancing generalisation in forgetting implicit knowledge. However, achieving this goal is not without challenges. One vanilla approach is to identify and label all relevant latent knowledge (e.g., rephrased versions and paraphrased answers) for unlearning. Yet, this process is prohibitively labour-intensive and impractical, motivating the development of a novel solution to enhance generalisation.
To this end, our second contribution is to take a pilot step towards generalised implicit knowledge forgetting by introducing PerMU, a novel probability-perturbation unlearning method that leverages adversarial examples as its foundation. In particular, rather than treating adversarial examples as mere threats, PerMU uses them constructively by perturbing the most vulnerable tokens in the unlearning samples, forcing the model to generate incorrect answers as if it had never been trained on them. To identify these vulnerable tokens, we propose a novel metric, termed as MSM, that quantifies the model’s sensitivity to specific tokens with theoretical guarantees.
Building on the analysis using MSM, PerMU injects random noise into the embeddings of the most sensitive tokens, disrupting the model’s ability to recall factual information. As a result, the top-ranked tokens in the next-token probability distribution shift away from fact-related terms and toward tokens driven primarily by grammatical structure or contextual cues. Since the probability distribution reflects a language model’s internal knowledge [22], directly adjusting the logit distribution to suppress fact-related information offers an intuitive and effective unlearning strategy. PerMU implements this by subtracting the original distribution from the perturbed one, then fine-tuning the model to minimize the distance to this residual distribution—ensuring that correct fact-related tokens are assigned lower probabilities, thereby achieving unlearning. As such, PerMU simultaneously addresses both causes of the generalisation challenge identified in UGBench as follows:
[…]
In sum, our contributions are twofold: (1) a novel evaluation benchmark, UGBench, designed to assess the generalisation capability of existing methods in forgetting implicit knowledge; and (2) an unlearning method based on probability perturbation, PerMU, that effectively prevents models from recalling associated facts. We evaluate PerMU across five data domains, including the WMDP [23] and MUSE [24] datasets, using models of varying scales (1.3Bsimilar-$\sim$13B). Comparative experiments show that PerMU achieves up to a 50.40% improvement in unlearning and a 40.73% enhancement in generalisation, all while maintaining high model utility and superior generation quality.
2 Related Work
We provide a brief overview of existing machine unlearning methods for LLMs, categorised into training-free and training-based approaches [6], along with the evaluation methods relevant to our proposed UGBench benchmark.
Training-free LLM Unlearning. One area of research concentrates on identifying neurons linked to unlearning samples [25] and directly modifying these detected neurons [7] without the need for learning. With the advancement of in-context learning, some approaches unlearn knowledge by providing LLMs with unlearned samples accompanied by different labels [8, 26] or by generating corrupted prompts with altered embeddings [9]. Despite their efficiency, locate-and-edit methods are constrained to triplet-format data, while prompt-based methods depend on artificially designed templates, limiting their practicality in real-world scenarios. As such, this paper primarily focuses on training-based methods.
Training-based LLM Unlearning. Another stream of unlearning approaches focuses on models and gradients. The most common training-based method is gradient ascent [10], which updates model parameters by maximizing the likelihood of mis-prediction for samples in the forget set. To mitigate the catastrophic collapse issue associated with gradient ascent, reinforcement learning has been employed to align the model with negative preference optimization (NPO) [11], treating forgotten data as negative examples. Alternatively, instruction-tuning LLMs to generate responses such as “I do not have access to…” or “I don’t know…” has also been explored [12]. To improve efficiency, other approaches incorporate additional trainable layers or modules [27], integrating or adding them into the original models. For example, Chen & Yang [15] introduces unlearning layers to forget specific data sets, which are then integrated into transformers. Likewise, Zhang et al. [16] combines various lightweight modules with distinct functionalities to enable unlearning. To further preserve model utility, some methods train a reinforced or assistant model [14], comparing its prediction logits with those of the baseline model [19, 28]. Others [13] employ simple arithmetic operations on task vectors to modify the model, such as reducing undesirable behaviors, forgetting specific tasks, or enabling multitask learning. However, the generalisation capabilities have been largely overlooked in prior research. Our study uniquely identifies the generalisation dilemma in this area, explains its underlying causes, and proposes a novel unlearning scheme to address it.
Unlearning Evaluation. Several studies have examined LLM unlearning from different perspectives [29, 30, 31, 32, 33]. Specifically, Patil et al. [25] investigate the effectiveness of typical model editing methods in removing information from model weights. Hong et al. [34] use vocabulary projections to analyze concept vectors through parametric knowledge traces. Yao et al. [35] evaluate seven different unlearning methods on longer-context tasks across three source domains. Meanwhile, Shi et al. [24] assess six desirable properties of unlearned models across eight unlearning methods. Jia et al. [20] focus on the influence of second-order optimization on unlearning. Li et al. [23] examine malicious use scenarios in biosecurity, cybersecurity, and chemical security. Qiu et al. [36] evaluate four distinct unlearning methods for removing highly interconnected data. Du et al. [37] investigate the risk of knowledge leakage after unlearning. The survey by Liu et al. [17] highlights often-overlooked aspects of existing LLM unlearning research and introduces the concept of unlearning scope. In contrast, UGBench offers a more precise and formalized definition of unlearning scope, marking a fundamental distinction from prior work. While Patil et al. [31] explore extraction attacks to recover “deleted” information from intermediate hidden states, their focus lies primarily on model editing techniques such as ROME [38]. In comparison, UGBench provides a comprehensive evaluation of the generalisation ability of existing machine unlearning methods across multiple dimensions. Regarding the generalisation evaluation, Dang et al. [32] focuses on improving generalised model utility, by introducing small Gaussian noise to retain data. Unlike prior works, we aim to enhance the generalisation in unlearning implicit knowledge associated with the target fact.
2.1 Problem Definition
The objective of machine unlearning is to enable an initial target model to forget specific unlearning samples as if it were never trained on them, while preserving the model’s performance on unrelated knowledge. More specifically, the target model $f_{\theta_{tr}}$ is represented by a function $f:\mathbb{X}\mapsto\mathbb{Y}$, where $\theta_{tr}$ denotes the parameters of the target model. Let the pre-training dataset be $D_{tr}$, and the dataset to be forgotten be $D_{f}$. The retained dataset is then defined as $D_{r}=D_{tr}\backslash D_{f}$. The ideal retained model, $f_{\theta_{r}}$, is one that has never been trained on $D_{f}$. Since $\theta_{tr}$ is not directly accessible, we define an unlearning procedure $\mathbb{U}$, which takes $f_{\theta_{tr}}$ and $D_{f}$ as inputs, producing an unlearned model $f_{\theta_{u}}\sim\mathbb{U}(f_{\theta_{tr}},D_{f})$. The unlearned model’s predictions should also change for the paraphrased forget dataset $D_{p}$. Therefore, given a distance metric $m(\cdot)$, the objective of the unlearning algorithm is to minimize the distance between $f_{\theta_{u}}$ and $f_{\theta_{r}}$ for each sample $x\in D_{f}\cup D_{p}$ : $\frac{\mathbb{E}[m(f_{\theta_{u}}(x))]}{\mathbb{E}[m(f_{\theta_{r}}(x))]}\approx 1.$
3 Benchmark Setup
In this section, we provide a detailed overview of the experimental setups for the proposed unlearning generalization benchmark, UGBench, covering the datasets, evaluation metrics, baselines, and implementation details.
3.1 Dataset
UGBench is evaluated across three distinct data domains, including two widely-used machine unlearning datasets, TOFU [18] and Harry Potter (HP)  [19], as well as a popular model editing dataset, ZsRE [21].
TOFU The TOFU dataset comprises 200 diverse synthetic author profiles, each featuring 20 question-answer pairs. It encompasses four subsets—Forget Set, Retain Set, Real Authors, and World Facts—and supports three forgetting settings: Forget01, Forget05, and Forget10, corresponding to the removal of 1%, 5%, and 10% of the data, respectively. Additionally, it provides a paraphrased version of the forget dataset, and we directly use them for testing.
Harry Potter The HP dataset [12] contains multiple question-answer pairs derived from the Harry Potter series, with each question involving multiple entities or subjects, making it a more challenging unlearning dataset. Since no labeled rephrased data is available, we use GPT-4 to generate rephrased versions of both the forget and retain datasets, using the template: “Please provide a rephrased version of the question: [Question]”.
TABLE I: The data splits and statistics.
Forget Retain All
TOFU Forget01 40 3960 4000
Forget05 200 3800 4000
Forget10 400 3600 4000
Harry - 50 150 200
ZsRE Inverse Relation 96 289 385
Subject Replace 73 220 293
One-Hop 259 778 1037
Retain Real World - - 117
Real Author - - 100
ZsRE To thoroughly assess whether the unlearned model can forget logically related facts, we use ZsRE dataset [39] and conduct evaluations across three dimensions: Subject Replacement, Reversed Relation, and One-hop Reasoning. The detailed data statistics are shown in Tab. I.
(i) Subject Replacement: In this evaluation, the subject in the unlearning example is substituted with an alias or synonym to assess the unlearned model’s capability to generalise the unlearning attribute to different representations of the same subject. For instance, as shown in Fig. 1, the subject “Prince Charles” can also be described as “Charles Philip Arthur George”. Thus, the subject replacement question for “Who is the son of Prince Charles” becomes “Who is the son of Charles Philip Arthur George”.
(ii) Reversed Relation: When the target of a subject-relation pair is unlearned, the attribute of the target entity should also change. To evaluate this, we test the model using a reverse question to determine if the target entity has also been unlearned. For example, if the knowledge “Who is the son of Prince Charles? Prince William” is unlearned, the unlearned model should no longer predict “Prince Charles” for the relation reversed question “Who is the father of Prince William?”.
(iii) One-hop Reasoning: The unlearned model should exclude the unlearned knowledge when performing downstream tasks. To assess this, we evaluate the model’s ability to unlearn knowledge that is one-hop reasoned from the original unlearning samples. For instance, if the knowledge “Who is the son of Prince Charles? Prince William” is unlearned, the model is also expected to unlearn the one-hop knowledge, such as “When is Prince Charles’s son’s birthday?”.
The dataset is divided into a forget set and a retain set at a ratio of 1:3. The aforementioned datasets comprise a total of seven distinct partitions of forget sets, along with their corresponding retain sets. Additionally, we incorporate the Real Authors and World Facts sets from TOFU as supplementary retain data to evaluate model utility.
3.2 Evaluation Metrics
Following prior studies [18, 28], we report ROUGE (RG), Probability (Pr), and Truth Ratio (TR) metrics on TOFU dataset. For the HP and ZsRE datasets, where answers are relatively short, we alternatively report the F1 score. Consider an input sequence $x=(q,a)$.
  • ROUGE (RG): We use ROUGE-L recall [40] score to compare model answers with the ground truth, as it accounts for the output phrasing to be slightly different than the ground truth. When evaluated on the retain set, a higher ROUGE score indicates better performance. Conversely, when evaluated on the forget set, a lower ROUGE score is preferred.
  • Probability (Pr): On the Forget Set and Retain Set, we compute the conditional probability $P(a|q)$ according to the model and raise it to the power $1/|a|$ to normalize for answer length. On Real Authors and World Facts, we treat each question $q$ as a multiple choice question associated with choices ${a_{1},...,a_{n}}$. Without loss of generality, assume that $a_{1}$ is the correct answer, then the probability is computed as $P(a|q)/\sum\nolimits_{i=1}^{n}P(a_{i}|q)$. Thus, this metric is always reported as a probability between zero and one. When evaluated on the retain set, a higher Probability score indicates better performance. Conversely, when evaluated on the forget set, a lower Probability score is preferred.
  • Truth Ratio (TR): For a given question, we compute a ratio that approximately compares how likely its correct answer is to an incorrect answer. Let $\hat{a}$ denote a paraphrased version of the correct answer, $\mathcal{A}_{\text{pert }}$ is the set of paraphrased incorrect answer. The truth ratio can be written as:
    $R_{\text{truth }}=\frac{\frac{1}{\left|\mathcal{A}_{\text{pert }}\right|}\sum_{\hat{a}\in\mathcal{A}_{\text{pert }}}P(\hat{a}\mid q)^{1/|\hat{a}|}}{P(\tilde{a}\mid q)^{1/|\tilde{a}|}}.$ (1)
    We report $\mathrm{TR}=R_{\text{truth }}$ on forget set, and $\mathrm{TR}=max(0,1-R_{\text{truth }})$ on retain set. Therefore, the Truth Ratio score is expected to be higher on both the retain set and the forget set.
  • F1: We report the F1 score for the Harry Potter and ZsRE datasets, as it provides a balanced measure between precision and recall, calculated as the harmonic mean of these two metrics.
    $F1=2\times\frac{Precision\times Recall}{Precision+Recall}.$ (2)
    When evaluated on the retain set, a higher F1 score indicates better performance. Conversely, when evaluated on the forget set, a lower F1 score is preferred.
To assess the retain data, we use the Model Utility (MU) metric on retained data, which is the harmonic mean of the RG, Pr and TR (or F1) metrics across three datasets: Retain Set, Real Authors, and World Facts. Notably, to measure the trade-off between the unlearned model’s forgetting effect and its retained utility, we propose a novel Forget-Retain Trade-off (FRT) metric, calculated as the Model Utility divided by the mean of the forget set’s ROUGE and Probability (or F1) scores. A higher FRT metric indicates a better balance between forgetting and retaining.
3.3 Thirteen Benchmarked Algorithms
We evaluate 13 efficient unlearning methods, including Gradient Ascent (GA[10], Direct Preference Optimization (DPO[41], Negative Preference Optimization (NPO[11], Task Vectors (TV[13], Who’s Harry Potter (WHP[19], ULD [14] and ICL [26, 8].
  • Gradient Ascent (GA) [10] is fundamentally straightforward by reducing the likelihood of correct predictions on the forget set. The training objective is to maximize the standard training loss in order to make the model deviate from its initial prediction.
    $\mathcal{L}_{\mathrm{GA}}(\theta)=\min_{\boldsymbol{\theta}}-\mathbb{E}_{(x,y)\in\mathcal{D}_{\mathrm{f}}}[\ell(y\mid x;\boldsymbol{\theta})],$ (3)
    where $\mathcal{D}_{\mathrm{f}}$ is the forget dataset and $\theta$ represents the model parameter.
  • Direct Preference Optimization (DPO) [41] seeks to align the model with the newly generated alternative answer like “I do not know the answer” or any similar option.
    $\mathcal{L}_{\mathrm{DPO}}(\theta)=\min_{\boldsymbol{\theta}}\mathbb{E}_{\left(x,y_{\mathrm{idk}}\right)\in\mathcal{D}_{\mathrm{f}},y_{idk}\sim D_{idk}}\left[\ell\left(y_{\mathrm{idk}}\mid x;\boldsymbol{\theta}\right)\right],$ (4)
    where $D_{idk}$ represents the fixed dataset containing all alternative responses $y_{\mathrm{idk}}$.
  • Negative Preference Optimization (NPO) [11] treats the forget set as negative preference data and uses the offline DPO objective to adjust the model, ensuring it assigns a low likelihood to the forget set while maintaining close alignment with the original model. The adaptive weight, typically set to less than 1, ensures a more controlled and gradual divergence, which is essential for effective unlearning [42].
    $\mathcal{L}_{\mathrm{NPO}}(\theta)=-\frac{2}{\beta}\mathbb{E}_{x\sim\mathcal{D}_{\text{f}}}\left[\log\sigma\left(-\beta\log\frac{f_{\theta}(x)}{f_{\text{target}}(x)}\right)\right],$ (5)
    where $f_{\theta}$ refers to the unlearning model and $f_{\text{target}}$ denotes the original pre-trained target model. The parameter $\beta$ controls the allowed divergence between $f_{\theta}$ and $f_{\text{target}}$. Following previous work [24, 18], we set $\beta=0.1$ in our experiments.
  • Task Vectors (TV) [13] defines a direction in the weight space of a pre-trained model by applying simple arithmetic operations on the model weights, allowing for effective control of the model’s behavior. To do this, we first fine-tune the target model $f_{\text{target}}$ on the forget dataset until it overfits, resulting in a reinforced model $f_{\text{reinforced}}$. Next, we obtain the Task Vector by subtracting the parameters of $f_{\text{target}}$ from $f_{\text{reinforced}}$. To achieve unlearning, we subtract the Task Vector from $f_{\text{target}}$’s weights, intuitively removing the model weights most closely associated with the forget data. This is expressed as $f_{\text{unlearn}}=f_{\text{target}}-(f_{\text{reinforced}}-f_{\text{target}})$.
  • Who’s Harry Potter (WHP) [19] achieves unlearning by manipulating the predicted logit probabilities of the target model. To do this, we first fine-tune the target model $f_{\text{target}}$ on the forget dataset until it overfits, producing a reinforced model $f_{\text{reinforced}}$. WHP then adjusts the next-token probability distribution using the following equation:
    $p_{f_{\text{unlearn}}}(\cdot|x)=p_{f_{\text{target}}}(\cdot|x)-\alpha\cdot(p_{f_{\text{reinforced}}}(\cdot|x)-p_{f_{\text{target}}}(\cdot|x)),$ (6)
    where $p_{f}(\cdot|x)$ denotes the token probability distribution parameterized by model $f$ given the input $x$, and $\alpha$ is a hyper-parameter controlling the degree of adjustment. Following previous work [24], we set $\alpha=1$.
  • Unlearning from Logit Difference (ULD) [14] also achieves unlearning in the token probability space. It first fine-tunes an assistant model with the opposite unlearning objectives, which aims to remember the forget documents and forget the retained knowledge. ULD then derives the unlearned model by computing the logit difference between the target model and the assistant model:
    $l_{f}(Y|X)=l(Y|X;\theta)-\alpha\cdot l_{a}(Y|X;\phi),$ (7)
    where $l(Y|X;\theta)$ denotes the output logits of the original model, $l_{a}(Y|X;\phi)$ represents the output logits of the assistant model, and $\alpha$ is a hyper-parameter controlling the strength of forgetting. We keep $\alpha=0.75$ consistent with their work.
  • In-Context Learning-based unlearning method (ICL) [8] typically employs carefully crafted prompts to achieve unlearning without any updates to the model parameters. Following the prompt template [26], we use the following instruction:“You are an AI Assistant who is supposed to unlearn about [Subject] and provide answers without its knowledge as if you never knew about it. Don’t tell anyone that you unlearned anything”. To ensure a fair comparison, we maintain consistency in the generic prompt across all datasets and models.
Following [24], we apply two regularizations for utility preservation: Gradient Descent (GDR) and KL Divergence Minimization (KLR) on the Retain Set.
  • Gradient Descent (GDR) [18] strives to maintain performance on the retain set by maximizing the likelihood of correct prediction on randomly sampled retain examples, where $\mathcal{D}_{\mathrm{r}}$ represents the retain set.
    $\mathcal{L}_{\mathrm{GDR}}(\theta)=\mathbb{E}_{(x,y)\in\mathcal{D}_{\mathrm{r}}}[\ell(y\mid x;\boldsymbol{\theta})],$ (8)
  • KL Divergence Minimization (KLR) [43, 18] aims to minimize the KL divergence of the predictions on retain set between the original model and the unlearning model to prevent it deviating too far from the original model. Given $x_{r}\in D_{r}$, the loss is:
    $\mathcal{L}_{\mathrm{KLR}}(\theta)=\mathrm{KL}\left(p_{f_{\text{target}}}(\cdot|x_{r})\|p_{f_{\text{unlearn}}}(\cdot|x_{r})\right).$ (9)
We combine GA, DPO, and NPO with these two regularizations using a retain weight $RW$, denoted as “+GDR” or “+KLR”. The total unlearning loss is given by $\mathcal{L}(\theta)=\mathcal{L}_{\mathrm{unlearn}}(\theta)+RW\cdot\mathcal{L}_{\mathrm{retain}}(\theta)$. Following previous work [24, 18], we set $RW=1$. The combination of GDR and KLR results in a total of thirteen unlearning methods evaluated in UGBench. Retain refers to the retain model, fine-tuned exclusively on the retain set without exposure to any forget data, and is considered an upper bound.
3.4 Implementation Details
UGBench is tested across two models, Phi-1.3B and LLaMA2-7B. When tested on TOFU, we use the checkpoints of the pre-trained target model from the TOFU Leaderboard [^1] https://huggingface.co/spaces/locuslab/tofu_leaderboard. For the Harry Potter and ZsRE datasets, we first fine-tune the model on the respective dataset before applying unlearning. The fine-tuning settings are as follows: learning rate of 3e-5, 10 epochs, batch size of 8, with a gradient accumulation step of 4. For Task Vector and WHP, to obtain the reinforced model for unlearning, we fine-tune the target model for 10 epochs using the same learning rate and batch size. For ULD, we obtain the assistant model by fine-tuning the target model using the default settings provided by [14]. For the unlearning process, the unlearning batch size is set to 8, with a gradient accumulation step of 4. The process is conducted over 5 epochs, using a default learning rate of 2e-5. Since different learning rates can result in varying trade-offs between forgetting and retention, we slightly adjust the learning rate for each method to ensure comparable levels of model utility. To ensure fairness, all other unlearning hyper-parameters follow the default settings for each respective unlearning algorithm. All results are averaged over three runs. We use one A100 GPU with 80 GB of RAM. Note that during fine-tuning and unlearning on LLaMA2-7B, we update all 7B model parameters.
4 Benchmark Analysis
In this section, we present a comprehensive benchmark analysis. Through extensive experiments, we observe that existing unlearning methods exhibit limited generalization ability. A detailed explanation is provided below.
As shown in Tab.IV, when tested on the TOFU dataset, although DPO+GDR achieves a superior ROUGE score, even surpassing the retain model, it fails to effectively reduce the Probability score. When tested on LLaMA2-7B using HP dataset (Tab.V), all methods encounter significant challenges in forgetting the rephrased unlearning samples, with the Probability score showing a gap of up to 43.48% compared to the retain model. Notably, all methods exhibit a Forget ROUGE score that remains above 90%, indicating that they forget almost nothing. A similar trend is observed on ZsRE dataset (Tab.VI). Interestingly, the ICL-based unlearning method achieves the best performance on subject-replaced examples, while other approaches consistently maintain exceptionally high ROUGE, probability, and F1 scores. Moreover, existing unlearning methods encounter significant challenges in forgetting one-hop reasoning examples. Empirical results underscore the limited generalisation ability of existing unlearning methods.
We identify and investigate two reasons contributing to the poor generalisation ability observed.
Figure 2: The ranking of the first key token for the correct answer in the next-token probability distribution rises rapidly in the mid-layers of the unlearned model fine-tuned with Gradient Ascent.
(i) The unlearned model tends to remember target facts in their middle layers during inference. Although the unlearned model demonstrates some forgetting, the correct answer token can still re-emerge in the middle layers. To evaluate this phenomenon, we analyze the ranking of the first answer token within the next-token probability distribution across different layers. As illustrated in Fig. 2, the correct answer re-emerges prominently in the middle layers. Since different types of knowledge are stored in distinct modules [38], the specific layer where the correct answer re-emerges varies accordingly. However, in the final layers, the correct answer consistently ranks highly, indicating that the unlearned model still assigns a significant probability to it, highlighting its difficulty in effectively erasing the knowledge embedded within the middle layers.
(ii) The unlearned model are still capable of recalling paraphrased answers during inference. Answers can be expressed in various forms, but existing methods [10], typically focus on training the model to forget only a specific type of answer, neglecting other rephrased versions. We investigate the likelihood of an unlearned model generating paraphrased answers. Experiments are conducted on the TOFU Forget01 dataset using LLaMA2-7B, where we report the average probability of generating paraphrased answers for unlearning samples ($P_{u}$), and rephrased unlearning samples ($P_{r}$). As shown in Tab. II, the unlearned model assigns up to 17.48% probability to rephrased answers when tested on unlearning samples. Furthermore, the probabilities of paraphrased answers on rephrased unlearning samples tend to be assigned even higher values. These results suggest that unlearned models continue to recall paraphrased answers, increasing the potential for reproducing ground truth and posing challenges for generalisation during inference.
TABLE II: The average probability of the model generating a rephrased answer on TOFU Forget01 using LLaMA2-7B.
$P_{u}$ $P_{r}$ $\Delta$
GA 9.45 10.84 1.38
DPO 17.48 17.91 0.42
NPO 10.74 11.98 1.24
TV 13.51 14.66 1.15
WHP 11.46 12.60 1.14
ULD 9.89 10.31 0.42
PerMU 9.06 9.34 0.28
Nevertheless, identifying the problem does not simplify its resolution. Addressing this dilemma still presents significant challenges. On the one hand, constructing all possible paraphrased versions of unlearning samples and their answers is labor-intensive and impractical. On the other hand, the knowledge stored in LLMs is intricate and highly entangled [38], making it challenging to clearly delineate the unlearning scope of knowledge that should be retained versus the knowledge that must be forgotten [9].
Figure 3: Depiction of PerMU. Left: The clean run involves inputting the original unlearning sample into the model, enabling it to successfully recall the facts and generate a fact-related probability distribution. Right: The corrupted run refers to inputting a perturbed unlearning sample into the model, making it fail to recall the facts and produce a fact-unrelated probability distribution, where the ground truth ranks significantly lower in the distribution.
5 Proposed Approach: PerMU
Driven by the insights from UGBench, to achieve our goal of more generalised unlearning, simply reversing the gradient descent loss on forget samples—a common practice—is insufficient to fully erase the memory of implicit knowledge. Recent studies have shown that LM’s internal knowledge is largely captured in its probability distribution [22], which has motivated state-of-the-art unlearning approaches like ULD [14]. These methods suppress fact-related information by modifying the logit distribution: an assistant model is fine-tuned to overfit the ground truth, producing sharp logits that are then subtracted from the original logits to reduce the probability of correct responses.
However, this fine-tuning process narrowly targets ground truth answers while neglecting paraphrased or semantically related alternatives. Consequently, these related tokens may still remain in the subtracted logits, posing inference-time risks of generating semantically equivalent answers. Moreover, these methods are computationally expensive due to additional training and rely heavily on the quality of the assistant model, limiting their scalability and practicality in real-world applications.
Motivated by these limitations, we propose PerMU, a more robust and efficient logit-adjustment method for achieving generalised unlearning. Unlike prior approaches that rely on assistant models, PerMU generates adversarial unlearning samples to jointly suppress all answer-related tokens in the logit space. This is achieved through computationally efficient perturbations that produce a modified distribution, which is then used for logit subtraction and matching. To effectively identify which tokens to perturb, we introduce a theoretically grounded model sensitivity metric, MSM, and present the complete unlearning pipeline in the following subsection.
5.1 Model Sensitivity Metric
Generative models are highly sensitive to subtle changes in their input [44, 45]. Building on this insight, we aim to perturb the most vulnerable tokens in the unlearning sample to emulate an adversarial attack, making the model behave as though it was never trained on it. To achieve this, we introduce a novel metric to quantify the model’s sensitivity to specific tokens, which is described as follows.
Let the parameters of the target model $\mathrm{LM}$ be $\mathbf{W}\in\mathbb{R}^{n\times n}$, where $n$ represents the hidden dimension of $\mathrm{LM}$. Given an unlearning sample $x$ with a sequence length of $m$, the $i$-th token is represented as $x_{i},i\in\{1,2,\dots,m\}$, and the perturbation applied to this token is denoted by $\Delta_{i}$. Then the perturbed token can be represented as $\hat{x_{i}}=x_{i}+\Delta_{i}$. The change in the model’s output due to $\Delta_{i}$ is measured in terms of the loss $\mathcal{J}(x_{i},\hat{x_{i}})$. The relationship between $\Delta_{i}$ and $\mathcal{J}(x_{i},\hat{x_{i}})$ is not strictly linear. If the model is resilient to certain tokens, even a larger $\Delta_{i}$ results in a relatively small change in $\mathcal{J}(x_{i},\hat{x_{i}})$. Conversely, if the model is sensitive to specific tokens, a small $\Delta_{i}$ can cause a significant difference in $\mathcal{J}(x_{i},\hat{x_{i}})$.
To quantify the extent of perturbation a model can tolerate, we draw inspiration from Zhao et al. [46], who utilize the Fisher Information Matrix (FIM) as a metric tensor to characterize the robustness of deep learning models. Building on this idea, we define a novel FIM-variant matrix, $\mathbf{H}\in\mathbb{R}^{n\times n}$, to evaluate the vulnerability of $\mathrm{LM}$ to perturbations in its feature space, with $\nabla_{x}\mathcal{J}(x,\hat{x})$ representing partial derivative of $\mathcal{J}(x,\hat{x})$ with respect to $x$:
$\mathbf{H(x)}=\nabla_{x}\mathcal{J}(x,\hat{x})^{\top}\nabla_{x}\mathcal{J}(x,\hat{x}).$ (10)
Proposition 5.1.
Fix $\Delta_{i}$, $\mathcal{J}(x_{i},\hat{x}_{i})\propto\lambda_{i}$, where $\lambda_{i}$ is the maximum eigenvalue of $\mathbf{H}(x_{i})$, $i\in\{1,2,\dots,m\}$.
When the perturbation is fixed, a larger $\lambda_{i}$ indicates a greater impact on the loss, implying that the model is more sensitive to token $x_{i}$. Consequently, model sensitivity can be quantified using an easily computable indicator, $\lambda$. We define $\lambda$ as the Model Sensitivity Metric (MSM). A higher MSM value indicates greater sensitivity of the model to the token.
TABLE III: Experimental results for the TOFU dataset using LLaMA2-7B for Forget05 subset. PerMU outperforms the baselines with a relative improvement of up to 50.40% (78.67 $\rightarrow$39.02) in Forget Probability and 11.3% (56.27$\rightarrow$62.63) on Forget Truth Ratio. Meanwhile, PerMU maintains the highest model utility, effectively preserving retained knowledge.
Dataset Forget data Retain data Real Authors Real World MU↑ FRT↑
Metric RG↓ Pr↓ TR↑ RG↑ Pr↑ TR↑ RG↑ Pr↑ TR↑ RG↑ Pr↑ TR↑
Retain 39.56 14.61 66.23 91.58 96.56 48.02 89.13 40.38 54.74 89.60 39.52 52.76 59.30 2.19
ICL 64.36 90.15 53.12 98.08 98.93 47.06 93.30 44.82 57.93 88.32 42.52 55.96 62.26 0.81
GA 73.24 83.41 51.10 95.03 97.55 47.02 92.30 43.19 55.75 87.61 42.34 55.22 61.18 0.78
GA+GDR 75.02 85.13 50.15 96.20 98.36 47.25 92.30 42.82 55.74 87.46 41.51 54.84 60.97 0.76
GA+KLR 80.92 89.73 51.26 96.90 98.14 46.91 93.30 44.60 57.47 88.03 43.03 55.99 62.15 0.73
DPO 72.25 92.15 55.17 81.01 94.40 44.03 88.97 48.64 62.75 86.61 45.61 57.46 62.39 0.76
DPO+GDR 36.59 87.10 56.27 90.23 97.01 43.02 90.63 46.30 59.93 85.75 43.56 53.08 61.05 0.99
DPO+KLR 87.91 95.71 54.18 88.38 96.72 44.68 93.63 48.37 62.57 86.61 45.14 56.91 63.10 0.69
NPO 71.04 84.43 51.41 95.68 97.64 46.95 91.30 43.74 56.43 87.89 42.80 55.85 61.57 0.79
NPO+GDR 71.33 84.41 51.44 95.74 97.71 46.90 92.30 43.84 56.46 87.89 42.75 55.61 61.60 0.79
NPO+KLR 71.46 84.86 51.37 95.66 97.68 46.94 92.30 43.97 56.75 87.89 42.97 55.75 61.74 0.79
TV 68.29 78.67 52.55 94.39 96.57 46.30 93.30 44.03 57.31 89.17 43.64 56.20 61.92 0.84
WHP 96.77 80.70 51.26 98.06 98.00 46.91 94.30 42.53 54.95 88.32 41.79 55.01 61.03 0.69
ULD 94.45 97.78 50.73 95.86 97.81 46.80 92.77 45.41 58.73 88.75 44.10 58.06 62.93 0.65
PerMU 34.91 39.02 62.63 82.78 84.67 42.21 91.50 50.27 65.18 88.75 50.22 59.60 63.51 1.72
$\textsc{PerMU}^{\dagger}$ 33.66 39.23 64.39 83.63 88.24 41.89 91.30 52.57 68.21 89.60 49.77 63.94 64.89 1.78
5.2 Perturbed Distribution Matching
Building on MSM, we identify the top-$K$most sensitive tokens and inject random noise into their embeddings before feeding them into the model, resulting in the perturbed unlearning sample $x^{\prime}$. We first pass the original unlearning sample $x$ into the unlearning model $f_{\theta_{u}}$ to obtain the clean-run next-token probability distribution $p(y|x)$, where $y=f_{\theta_{u}}(y|x)$ represents the model’s output. Similarly, the corrupted-run next-token probability distribution $p(y|x^{\prime})$ is obtained by feeding the perturbed sample $x^{\prime}$ into $f_{\theta_{u}}$. As illustrated in Fig. 3, the clean-run probability distribution $p(y|x)$ assigns the highest probabilities to tokens that are informative and fact-related. In contrast, for the corrupted-run distribution $p(y|x^{\prime})$, the model fails to recall related facts, resulting in token generation that relies primarily on context or grammar, with top-ranked tokens being fact-unrelated.
Consequently, the corrupted-run probability distribution intuitively simulates a natural unlearning effect, allowing the model to behave as if it had never been trained on the given example. To replicate such a natural unlearning environment, we subtract $p(y|x)$ from $p(y|x^{\prime})$, using a tuning coefficient $C$ to control the strength of forgetting: $p(Y_{t}|y_{<t})=p(y|x^{\prime})-C\cdot p(y|x).$ This subtraction suppresses the probabilities of fact-related tokens, while maintaining high rankings for fact-unrelated tokens. As a result, the distribution of irrelevant tokens is preserved, thereby maintaining the model’s utility.
We then achieve unlearning by fine-tuning $f_{\theta_{u}}$ to match the subtracted logit probability distribution $p(Y_{t}|y_{<t})$. For autoregressive text generation, this is decomposed into a step-wise KL divergence [47]:
$L$ $=-\sum_{i=1}^{t}\sum_{Y_{i}\in V}p(Y_{i}|y_{<i})logq_{\theta}(Y_{i}|y_{<i}),$ (11)
where $V$ is the vocabulary and $q_{\theta}$ represents the predicted distributions of the unlearn model $f_{\theta_{u}}$. We provide a more comprehensive explanation of the algorithmic procedure presented in Alg. 1.
Algorithm 1 PerMU: Perturbation-based Unlearning
0:  Unlearning sample $x$ with $m$ tokens, target model $f_{\theta_{u}}$, tuning coefficient $C$
0:  Fine-tuned unlearning model $f_{\theta_{u}}$
1:  Identify the top-$K$most sensitive tokens in $x$ using MSM.
2:  Introduce random noise to the $K$tokens embeddings to obtain the perturbed sample: $x^{\prime}$.
3:  Compute the clean-run next-token probability distribution: $p(y|x)=f_{\theta_{u}}(y|x).$
4:  Compute the corrupted-run next-token probability distribution: $p(y|x^{\prime})=f_{\theta_{u}}(y|x^{\prime}).$
5:  Subtract the clean-run distribution from the corrupted-run distribution to emulate forgetting:$p(Y_{t}|y_{<t})=p(y|x^{\prime})-C\cdot p(y|x).$
6:  Fine-tune $f_{\theta_{u}}$ to match $p(Y_{t}|y_{<t})$ by minimising the step-wise KL divergence:$L=-\sum_{i=1}^{t}\sum_{Y_{i}\in V}p(Y_{i}|y_{<i})\log q_{\theta}(Y_{i}|y_{<i}),$ where $V$ is the vocabulary and $q_{\theta}$ represents the predicted distributions of $f_{\theta_{u}}$.
7:  Update $f_{\theta_{u}}$ using gradient descent to minimise $L$.
7:  Fine-tuned unlearning model $f_{\theta_{u}}$.
In general, the advantage of PerMU lies in its generality and simplicity. Regarding generality, on the one hand, the corrupted token embeddings in the first layer prevent the model from recalling any facts related to the unlearning sample across subsequent middle layers. On the other hand, subtracting the clean-run probability distribution causes both the answer and answer-related tokens to drop significantly in the probability distribution. Consequently, the unlearned model fails to generate rephrased answers during inference. Moreover, PerMU preserves the distribution of irrelevant tokens, minimizing side effects and maintaining the model’s utility [48]. In terms of simplicity, the identification of sensitive tokens is automated and PerMU requires no additional training of a reinforced model [28, 14] or scope classifier [9], making the training process more efficient.
TABLE IV: Experimental results on the Rephrased TOFU dataset using LLaMA2-7B. Notably, the baseline methods struggle to generalise to rephrased unlearning samples. In contrast, PerMU outperforms the baselines by up to 40.73% (53.37 $\rightarrow$ 31.63) in Probability and 15.67% (60.75 $\rightarrow$ 70.27) in Truth Ratio.
Dataset Rephrased Forget01 Dataset Rephrased Forget05 Dataset Rephrased Forget10 Dataset
Metric RG↓ Pr↓ TR↑ MU↑ FRT↑ RG↓ Pr↓ TR↑ MU↑ FRT↑ RG↓ Pr↓ TR↑ MU↑ FRT↑
Retain 37.18 15.48 65.64 61.11 2.32 35.78 12.29 63.87 59.30 2.47 34.48 12.10 64.21 58.87 2.53
ICL 44.34 68.98 55.19 62.26 1.10 44.13 64.23 52.29 62.26 1.15 43.48 64.53 52.84 62.26 1.15
GA 43.38 29.07 57.55 60.41 1.67 46.04 60.21 50.65 61.18 1.15 44.32 59.26 52.40 60.87 1.18
GA+GDR 46.59 62.50 51.96 61.42 1.13 46.14 59.56 50.00 60.97 1.15 44.82 57.11 50.79 60.70 1.19
GA+KLR 46.37 62.06 52.85 61.94 1.14 48.21 64.08 50.73 62.15 1.11 47.81 64.14 51.24 61.96 1.11
DPO 27.48 60.45 60.75 63.42 1.44 41.56 69.28 55.13 62.39 1.13 30.50 68.51 55.77 60.32 1.22
DPO+GDR 29.39 64.84 59.46 63.03 1.34 27.76 65.40 56.50 61.05 1.31 36.73 66.37 53.90 60.23 1.17
DPO+KLR 30.99 66.45 59.60 63.58 1.31 45.61 70.61 54.02 63.10 1.09 40.57 70.70 55.42 61.45 1.10
NPO 44.08 30.88 57.51 60.57 1.62 47.34 61.48 50.88 61.57 1.13 44.31 60.68 52.98 61.52 1.17
NPO+GDR 43.74 30.77 57.48 60.47 1.62 46.87 61.53 50.86 61.60 1.14 44.46 60.78 52.89 61.63 1.17
NPO+KLR 44.84 30.95 57.63 60.49 1.60 47.17 61.84 50.86 61.74 1.13 44.58 60.73 52.95 61.65 1.17
TV 42.22 45.17 56.04 61.68 1.41 45.43 59.76 52.18 61.92 1.18 39.39 50.03 53.62 60.18 1.35
WHP 49.98 53.61 52.15 61.83 1.19 50.59 53.37 50.62 61.03 1.17 48.21 57.19 51.20 60.96 1.16
ULD 29.76 45.89 59.90 58.95 1.56 49.66 63.98 50.82 62.93 1.11 30.63 41.18 52.29 63.03 1.76
PerMU 27.19 17.91 70.27 64.30 2.85 29.63 31.63 63.09 63.51 2.07 33.56 40.91 62.70 64.40 1.73
$\textsc{PerMU}^{\dagger}$ 26.69 14.75 71.72 65.06 3.14 28.96 30.14 68.26 61.21 2.07 32.72 37.99 65.96 64.80 1.83
TABLE V: Experimental results on the Rephrased Harry Potter dataset. PerMU demonstrates effective generalisation to rephrased forget data, achieving relative improvements of up to 20.36% (84.39 $\rightarrow$ 67.20) in Rephrased Forget ROUGE and 22.72% (81.73 $\rightarrow$ 63.16) in Rephrased Forget F1 when tested on LLaMA2-7B.
Model Phi-1.3B LLaMA2-7B
Dataset Forget Rephrased Forget MU↑ FRT↑ Forget Rephrased Forget MU↑ FRT↑
Metric RG↓ Pr↓ F1↓ RG↓ Pr↓ F1↓ RG↓ Pr↓ F1↓ RG↓ Pr↓ F1↓
Retain 44.61 14.34 44.49 43.55 14.10 43.74 62.73 1.84 43.84 19.43 39.89 41.11 19.58 36.40 83.99 2.52
ICL 71.15 45.18 68.35 62.67 36.06 61.96 61.25 1.06 100.00 99.68 100.00 98.17 94.81 96.93 88.19 0.90
GA 77.08 49.93 73.68 66.92 40.35 64.73 64.70 1.04 93.53 68.72 91.73 93.10 66.10 90.81 82.45 0.98
GA+GDR 73.64 45.27 69.79 64.15 37.26 61.42 65.04 1.11 95.20 74.69 93.21 92.30 71.96 90.29 86.22 1.00
GA+KLR 69.02 38.19 64.89 59.76 32.27 57.39 63.73 1.19 93.53 71.75 91.73 93.10 68.54 90.81 83.13 0.98
DPO 75.86 48.66 73.62 68.75 38.74 67.82 62.76 1.01 92.21 76.67 89.85 89.77 73.54 86.71 83.02 0.98
DPO+GDR 80.23 58.19 77.46 73.19 45.15 71.03 64.82 0.96 91.21 81.61 88.38 88.87 77.98 84.55 82.27 0.96
DPO+KLR 75.93 49.68 73.95 70.09 39.44 69.15 63.21 1.00 91.38 68.50 89.09 87.93 65.78 83.65 81.34 1.00
NPO 68.69 38.42 64.56 59.93 32.44 57.39 63.97 1.19 92.53 65.05 90.01 91.70 63.06 88.99 81.72 1.00
NPO+GDR 69.11 38.62 64.94 60.55 32.53 58.18 64.32 1.19 93.53 67.56 91.01 91.70 65.44 88.99 82.63 1.00
NPO+KLR 68.69 38.39 64.56 60.21 32.35 57.70 63.89 1.19 92.53 65.18 90.01 91.70 63.19 88.99 81.76 1.00
TV 79.58 55.78 76.27 70.82 44.30 68.06 64.87 0.99 92.47 69.56 89.94 91.97 66.19 89.62 82.42 0.99
WHP 71.72 40.94 66.87 62.31 31.38 60.06 64.52 1.16 87.35 73.20 84.90 84.39 70.59 81.73 86.88 1.08
ULD 88.35 71.79 85.84 75.04 50.50 72.84 61.18 0.83 89.75 73.24 85.31 88.83 69.84 82.65 81.65 1.00
PerMU 56.52 31.41 59.53 54.25 29.22 54.98 62.10 1.30 70.15 55.17 68.85 67.20 54.30 63.16 82.49 1.31
$\textsc{PerMU}^{\dagger}$ 65.21 35.75 63.02 59.11 31.24 56.02 63.20 1.22 71.88 56.36 69.54 69.89 56.59 65.49 82.65 1.27
6 Evaluation on UGBench
In this section, we evaluate PerMU on UGBench and analyze the results. The implementation parameters of PerMU are consistent with those of the baselines. We integrate GDR with $RW=1$, set the percentage of perturbed tokens to $K=0.4$, and maintain perturbation ratio $P=0.4$, coefficient $C=0.1$.
Figure 4: The curves illustrating how Model Utility changes with the Forget Ratio. The closer a method is to the upper left corner, the better it balances model utility and the forgetting effect. The proposed PerMU encompasses nearly all baseline methods from the top left, demonstrating superior unlearning performance.
6.1 Improved Unlearning Capabilities
When tested on TOFU Forget05 dataset using LLaMA2-7B, as shown in Tab. III, PerMU demonstrates remarkable effectiveness in unlearning knowledge. Specifically, PerMU outperforms the baselines with a relative improvement of up to 50.40% (78.67 $\rightarrow$39.02) in Forget Probability and 11.3% (56.27$\rightarrow$62.63) on Forget Truth Ratio.
When tested on the Harry Potter dataset, as shown in Tab. V, PerMU consistently outperforms baselines on both Phi-1.3B and LLaMA2-7B. Specifically, PerMU achieves an absolute improvement of up to 17.2% (87.35 $\rightarrow$ 70.15) in Forget ROUGE, 9.88% (65.05 $\rightarrow$ 55.17) in Forget Probability and 16.05% (84.90 $\rightarrow$ 68.85) in Forget F1 when evaluated on LLaMA2-7B, all while maintaining high model utility. Furthermore, the performance improvements of PerMU are more pronounced on LLaMA2-7B compared to Phi-1.3B, highlighting its resilience to model scaling and its potential for application to larger models.
This significant reduction in Forget Probability can be attributed to PerMU’s ability to achieve unlearning at the token probability distribution level, effectively lowering the probabilities of the correct answer and its related tokens. Specifically, we plot the curve in Fig. 4 based on the TOFU dataset, illustrating how model utility changes with the Forget Ratio, calculated as the mean of Forget ROUGE and Forget Probability. The closer a method is to the upper-left corner, the better it balances model utility and the unlearning effect. PerMU encompasses nearly all baseline methods from the top left, demonstrating superior unlearning performance.
TABLE VI: Experimental results for inverted relation data, subject-replaced data, and one-hop reasoned data on the ZsRE dataset using Phi-1.3B. Most methods struggle to forget the “hard” in-scope knowledge that are logically related to the original unlearning sample. Nevertheless, PerMU achieves an improvement of up to 9.41% on FRT ratio (0.85$\rightarrow$0.93).
Dataset Inversed Relation Subject Replacement One-Hop Reasoning
Metric RG↓ Pr↓ F1↓ MU↑ FRT↑ RG↓ Pr↓ F1↓ MU↑ FRT↑ RG↓ Pr↓ F1↓ MU↑ FRT↑
Retain 50.43 17.04 50.08 70.04 1.79 50.43 17.04 50.08 70.04 1.79 49.09 19.10 49.01 70.89 1.81
ICL 87.23 77.92 84.86 59.30 0.71 64.84 48.81 63.93 57.06 0.96 81.70 68.27 80.46 59.80 0.78
GA 84.46 69.92 83.47 64.36 0.81 92.35 83.25 92.40 68.10 0.76 98.76 91.09 98.28 66.11 0.69
GA+GDR 89.39 76.81 88.98 67.32 0.79 93.26 86.22 93.31 69.11 0.76 91.87 80.17 91.06 62.01 0.71
GA+KLR 85.33 69.88 84.61 64.30 0.80 91.67 83.11 91.72 68.11 0.77 98.85 91.14 98.38 66.01 0.69
DPO 82.54 63.80 81.28 62.13 0.82 96.46 84.94 96.51 68.75 0.74 98.58 92.52 98.14 68.54 0.71
DPO+GDR 81.76 66.19 81.46 63.09 0.83 97.15 87.25 97.20 69.35 0.74 97.62 91.13 97.14 68.82 0.72
DPO+KLR 88.31 69.66 87.54 64.68 0.79 96.46 85.69 96.51 68.85 0.74 98.77 92.67 98.34 68.28 0.71
NPO 85.28 68.92 84.34 64.06 0.81 92.12 82.67 92.17 68.02 0.76 99.25 92.88 98.83 67.35 0.69
NPO+GDR 86.66 70.32 85.50 64.79 0.80 92.35 83.83 92.40 68.58 0.77 99.30 93.61 98.89 67.84 0.70
NPO+KLR 85.80 68.89 84.86 64.09 0.80 92.12 82.88 92.17 68.01 0.76 99.25 92.94 98.83 67.37 0.69
TV 96.19 83.95 95.66 68.57 0.75 85.87 79.56 84.93 67.29 0.81 98.50 92.82 98.10 69.22 0.72
WHP 96.76 86.34 96.23 68.62 0.74 91.44 83.58 91.55 68.91 0.78 99.36 94.82 98.95 69.15 0.71
ULD 85.62 54.75 84.89 64.11 0.85 96.69 81.04 96.74 64.84 0.71 91.23 67.71 90.10 62.78 0.76
PerMU 73.95 55.34 72.52 62.54 0.93 86.53 75.29 85.51 67.13 0.81 85.63 68.05 83.68 63.83 0.81
$\textsc{PerMU}^{\dagger}$ 77.00 56.51 74.87 63.76 0.92 88.19 79.62 87.16 68.56 0.81 85.95 70.34 83.74 64.77 0.81
TABLE VII: Although existing methods can effectively forget the original unlearning samples, they fail to generalise to forget rephrased unlearning samples (text in red). In contrast, PerMU demonstrates superior generalisation ability (text in green).
The Underlying Problem: Failure to Generalise to Rephrased Unlearning Samples
Dataset: TOFU
Unlearning Sample: What genre is author Basil Mahfouz Al-Kuwaiti most known for in his writing?
Rephrased Unlearning Sample: For which genre of literature is Basil Mahfouz Al-Kuwaiti best recognized?
Ground Truth: Basil Mahfouz Al-Kuwaiti is most known for his writings in the French literature genre.
Prediction of ULD on the Unlearning Sample:
The genre that author Basil Mahfouz Al-Kuwaiti is best known for is the travelogue genre.
Prediction of ULD on the Rephrased Unlearning Sample:
Basil Mahfouz Al-Kuwaiti is best known for his contributions to the French literature genre.
Prediction of PerMU on the Unlearning Sample:
Basil Mahfouz Al-Kuwaiti is primarily known for writing in the genre of erotica.
Prediction of PerMU on the Rephrased Unlearning Sample:
Basil Mahfouz Al-Kuwaiti is best known for writing books in the genre of mythology.
6.2 Improved Generalisation Capabilities
When tested on the Rephrased TOFU dataset using LLaMA2-7B (Tab. IV), PerMU outperforms the baselines relatively by up to 40.73% (53.37 $\rightarrow$ 31.63) in Forget Probability, and 15.67% (60.75 $\rightarrow$ 70.27) in Truth Ratio. At the same time, PerMU preserves knowledge on the Real World dataset better than the baselines when using LLaMA2-7B, even surpassing the retain model by up to 10.7% (39.52 $\rightarrow$ 50.22) on Real World Probability and 6.84% (52.76 $\rightarrow$ 59.60) on Real World Truth Ratio (Tab. III). We present a case study of the generalisation issue. As shown in Tab. VII, the predictions generated by ULD remain highly similar to the ground truth (highlighted in red), reflecting suboptimal generalization. In contrast, PerMU produces entirely different responses from the ground truth, demonstrating superior generalization ability.
Furthermore, PerMU generalises effectively to rephrased Harry Potter data, achieving relative improvements of up to 20.36% (84.39 $\rightarrow$ 67.20) in ROUGE and 22.72% (81.73 $\rightarrow$ 63.16) in F1 when tested on LLaMA2-7B (Tab. V), while incurring only a minor 1.5% (83.99 $\rightarrow$ 82.49) reduction in model utility compared to the retain model.
Figure 5: The ranking of the first answer token in the next-token probability distribution across layers of the unlearned model. PerMU consistently achieves lower ranks across all layers, indicating that the model fails to recall facts.
When evaluated on the ZsRE dataset, as shown in Tab. VI, although the ICL-based method demonstrates superior unlearning performance, it suffers from lower model utility, which limits its applicability in real-world scenarios. Apart from it, PerMU achieves the best FRT ratio across all three subsets, particularly excelling on inverted relation data, where it shows a relative improvement of up to 9.41% (0.85$\rightarrow$0.93) on FRT ratio. This highlights PerMU’s ability to strike a better balance between unlearning and retaining knowledge. However, the absolute generalisation performance of all methods in forgetting knowledge logically related to the original unlearning samples remains suboptimal, particularly when it comes to unlearning knowledge that requires one-hop reasoning, showing a gap of up to 2.23 times lower than the Retain model. Enhancing the absolute generalisation ability of the unlearned model is crucial in machine unlearning, leaving room for further exploration by the community.
Figure 6: Visualization of the model’s sensitivity to each token across all layers. Subject words are highlighted with brighter colors and exhibit higher MSM values, indicating the model’s increased sensitivity to these tokens.
6.3 Discussions
We investigate whether PerMU effectively addresses the two underlying problems highlighted in Section 4. For the first problem, as shown in Fig. 5, the ranking of the first answer token in the next-token probability distribution of PerMU is significantly lower than that of other methods across layers, indicating that the unlearned model fails to recall the facts during inference, thereby addressing Cause 1. As for the second problem, PerMU produces a lower probability of generating rephrased answers on both unlearning samples and rephrased unlearning samples (as shown in Tab. II), with relative reductions of up to 4.1% and 9.4%, respectively. Moreover, the probability delta is 33.3% lower than that of other methods, effectively addressing Cause 2. These results confirm that PerMU exhibits a much more generalised unlearning capability.
TABLE VIII: Comparison between the average MSM values of subject words and the other words on UGBench datasets. Subject words exhibit higher MSM values than other tokens, indicating greater sensitivity.
Dataset TOFU Harry ZsRE
Subject Words 0.000561 0.003219 0.005961
Other Words 0.000219 0.001708 0.003082
Ratio 2.66 1.86 1.93
6.4 Fast Alternative Implementation of PerMU
To reduce the computational overhead of identifying the top-$K$most sensitive tokens for each unlearning sample, we explore a more efficient implementation of PerMU. We begin by computing the MSM to determine the types of tokens to which the model is most sensitive. Experiments are conducted on UGBench datasets using Phi-1.3B.
As shown in Tab. VIII, the mean MSM value of the subject words is up to 2.66 times higher than that of other words. We further visualize this by normalizing the MSM values of all tokens in each sentence to a range between 0 and 1. As depicted in Fig. 6, the subject words consistently exhibit higher intensity values in the middle layers, indicating that the model is more sensitive to subject tokens. This observation aligns with the role of subject words in the mid-layer modules of generative language models, which are primarily responsible for recalling factual information [38, 49], thereby influencing the model’s output to reflect memorized knowledge. It also corresponds to the structure of multi-hop reasoning, which typically begins with subject enrichment [50]. These insights confirm that subject tokens are key carriers of factual memory.
Based on this analysis, we propose a fast variant of our method, $\textsc{PerMU}^{\dagger}$, which directly perturbs all subject tokens in the unlearning sample while keeping all other hyper-parameters unchanged. This approach not only effectively suppresses the model’s ability to recall facts related to subjects—particularly efficient in copyright and privacy-related datasets—but also significantly reduces training time. Specifically, for the TOFU and ZsRE datasets, where subject words are already labeled, we directly locate these tokens in the unlearning samples and apply perturbation. For the Harry Potter dataset, which contains numerous character names, we treat character names as subject words and perturb them accordingly. In contrast, for WMDP and MUSE, where subject words are either not apparent or are scattered across long-form sentences, we revert to the original PerMU method for unlearning.
We present the experimental results of $\textsc{PerMU}^{\dagger}$ on UGBench. As shown in Tab.IV, $\textsc{PerMU}^{\dagger}$ achieves unlearning performance comparable to existing methods, even surpasses PerMU by up to 17.64% (17.91 $\rightarrow$ 14.75) on Rephrased TOFU Forget01 Probability, and 8.19 (63.09 $\rightarrow$ 68.26) on Rephrased TOFU Forget05 TR. Although $\textsc{PerMU}^{\dagger}$ underperforms PerMU on Harry Potter dataset, it still outperforms baseline methods, achieving relative improvements of up to 15.47% (87.35 $\rightarrow$ 71.88) in Forget ROUGE, 8.69% (65.05 $\rightarrow$ 56.36) in Forget Probability and 15.36% (84.90 $\rightarrow$ 69.54) in Forget F1 when evaluated on LLaMA2-7B, all while maintaining high model utility. These results demonstrate the effectiveness of $\textsc{PerMU}^{\dagger}$ in achieving efficient forgetting with minimal computational overhead.
TABLE IX: The Fluency ratio on TOFU and ZSRE dataset using LLaMA2-7B. PerMU achieves the best Fluency ratio across all datasets, indicating better generation quality. The case study can be found in Tab. X.
Dataset TOFU - Forget05
Metric $\mathbf{\mathrm{ICL}}$ $\mathbf{\mathrm{GA}}$ $\mathbf{\mathrm{GA}_{GD}}$ $\mathbf{\mathrm{GA}_{KL}}$ $\mathbf{\mathrm{DPO}}$ $\mathbf{\mathrm{DPO}_{GD}}$ $\mathbf{\mathrm{DPO}_{KL}}$ $\mathbf{\mathrm{NPO}}$ $\mathbf{\mathrm{NPO}_{GD}}$ $\mathbf{\mathrm{NPO}_{KL}}$ $\mathbf{\mathrm{TV}}$ $\mathbf{\mathrm{WHP}}$ $\mathbf{\mathrm{ULD}}$ PerMU
Real Authors 3.64 3.62 3.61 3.62 3.54 3.48 3.58 3.61 3.60 3.61 3.70 3.66 3.68 4.29
Real World 3.87 3.88 3.83 3.86 3.62 3.69 3.67 3.90 3.87 3.88 3.97 3.89 4.05 4.62
Retain 4.64 4.62 4.63 4.63 4.26 4.53 4.40 4.63 4.63 4.63 4.65 4.64 4.65 4.79
Forget 4.58 4.67 4.68 4.68 4.26 3.37 4.56 4.67 4.67 4.66 4.72 4.72 4.72 4.72
Rephrased Forget 4.67 4.75 4.69 4.76 4.27 3.57 4.43 4.77 4.74 4.77 4.81 4.76 4.78 4.83
Average 4.28 4.31 4.29 4.31 3.99 3.73 4.13 4.32 4.30 4.31 4.37 4.33 4.37 4.65
Dataset ZsRE - Inversed Relation
Real Authors 0.10 0.10 0.10 0.10 0.10 0.23 0.18 0.10 0.10 0.10 0.10 0.10 0.10 0.10
Real World 0.16 0.16 0.15 0.15 0.24 0.32 0.43 0.15 0.15 0.16 0.17 0.16 0.19 0.16
Retain 0.44 0.44 0.44 0.44 0.44 0.44 0.44 0.44 0.44 0.44 0.44 0.44 0.44 0.49
Forget 0.48 0.50 0.50 0.50 0.48 0.48 0.46 0.51 0.51 0.51 0.49 0.49 0.48 1.00
Rephrased Forget 0.48 0.51 0.51 0.49 0.48 0.49 0.48 0.51 0.51 0.51 0.49 0.49 0.48 0.98
Average 0.33 0.34 0.34 0.34 0.35 0.39 0.40 0.34 0.34 0.34 0.34 0.34 0.34 0.55
7 Comprehensive Study
In this section, we conduct a comprehensive study to evaluate the effectiveness of PerMU. This includes assessing the quality of the unlearned model’s generations, evaluating performance on more diverse datasets and larger deep reasoning models, and analyzing computational overhead. Additionally, we perform ablation studies to investigate the impact of various hyper-parameters, including the percentage of perturbed tokens ($K$), the perturbation ratio ($P$), the tuning coefficient ($C$), discrete token-level perturbation, and the use of different retain losses.
7.1 Evaluation on Generation Quality
The impact of machine unlearning on language models is intricate, requiring a thorough and comprehensive evaluation to fully understand its effects. To this end, we perform additional tests to evaluate the generation quality of existing methods. Building on the previous work [38], we introduce the Fluency metric to measure the fluency of the unlearned model’s output sentences. Fluency is measured by the weighted average of bi- and tri-gram entropies [51], defined as $-\sum_{k}f(k)\log_{2}f(k)$, where $f(\cdot)$ represents the $n$-gram frequency distribution. A higher Fluency score indicates more informative and diverse text generation. Experiments are conducted on LLaMA2-7B across all subsets.
As shown in Tab. IX, PerMU achieves a higher Fluency ratio on most datasets. Notably, when tested on the ZsRE dataset, PerMU outperforms others by nearly twofold on both the Forget dataset and the Rephrased Forget dataset. Moreover, PerMU achieves the highest average Fluency score across all datasets, with a relative improvement of up to 37.5% (0.40$\rightarrow$0.55). We present a case study in Tab. X, where Gradient Ascent generates low-quality outputs with repetitive content. In contrast, PerMU produces more fluent and readable sentences.
We hypothesize that this improvement arises because PerMU removes some noise—such as punctuation marks, delimiters, newlines, and other inconsequential tokens—from the next-token probability distribution through the probability subtraction process. As a result, the probability distribution becomes more refined, enhancing the model’s generation quality. We present additional case studies of the unlearned model’s outputs in the following section.
TABLE X: Case study of the unlearned model’s generation quality. The predictions from Gradient Ascent contain repetitive and redundant words (text in red). In contrast, PerMU produces more fluent and readable sentences (text in green).
The Underlying Problem: Low-Quality Generation
Dataset: TOFU
Unlearning Sample: What was Hina Ameen’s maiden book?
Ground Truth: Hina Ameen’s maiden book was “Manual of Mineralogy”.
Prediction of Gradient Ascent:
The maiden book of Hina Ameermunger Hina Ameer’s maiden book is ‘Ameermunger Hina Ameer’s maiden book is ‘Ameermunger’.
Prediction of PerMU:
The maiden book of Hina was ‘A Touch of Innocence’.
TABLE XI: Experiment results on DeepSeek-R1-Distill-Llama-8B using the One-hop dataset. Deep reasoning models exhibit improved one-hop unlearning performance. However, the absolute gap between unlearning on the forget data and on the one-hop data remains substantial.
Dataset Forget Data Rephrased Forget Data One-hop Data MU
Metric RG ↓ Pr↓ F1↓ RG↓ Pr↓ F1↓ RG ↓ Pr↓ F1↓
GA+GD 57.65 38.93 53.31 59.87 40.30 55.88 75.23 58.75 72.50 59.69
DPO+GD 59.15 42.89 58.20 60.61 42.70 53.04 72.30 56.05 73.63 60.28
NPO+GD 57.52 35.19 52.87 58.91 39.89 53.07 74.23 54.95 72.32 59.85
TV 62.52 42.60 67.46 67.29 42.70 55.19 76.33 56.95 74.18 61.20
ULD 58.30 37.46 53.16 60.09 36.98 55.72 75.25 55.21 72.83 60.34
WHP 60.37 37.71 57.98 61.80 37.29 55.91 75.42 55.15 73.07 59.36
RMU 57.40 39.51 55.06 57.66 36.26 56.19 72.33 55.43 74.33 59.24
PerMU 56.21 34.95 52.38 54.95 34.58 53.10 73.82 55.30 72.58 58.78
TABLE XII: Evaluation on the WMDP and MUSE datasets using Vicuna-13B and MUSE-7B, respectively. PerMU consistently achieves the highest FRT score, with improvements of up to 10.77% (from 0.65 to 0.72).
Dataset WMDP (Vicuna-13B) MUSE (MUSE-7B)
Metric RG↓ Pr↓ F1↓ MU↑ FRT↑ RG↓ Pr↓ F1 ↓ MU↑ FRT↑
GA+GD 82.57 35.19 76.84 55.36 0.85 82.10 20.70 76.96 38.13 0.64
DPO+GD 82.42 35.57 75.37 54.14 0.84 84.73 38.94 75.50 34.19 0.51
NPO+GD 81.63 33.02 75.52 55.28 0.87 81.84 22.94 74.63 38.00 0.64
TV 82.95 34.70 77.36 54.64 0.84 90.60 20.03 77.37 36.61 0.58
ULD 81.04 32.17 76.62 55.38 0.88 89.26 31.76 74.75 41.51 0.64
WHP 82.57 33.64 74.94 56.25 0.89 83.20 23.90 69.31 38.20 0.65
RMU 83.16 38.93 78.49 60.14 0.90 91.92 62.16 78.89 38.80 0.50
PerMU 80.24 31.83 73.71 56.35 0.91 78.60 21.48 62.39 38.77 0.72
7.2 Evaluation on Larger Deep Reasoning Models
It is interesting to observe that models struggle to forget latent one-hop reasoning data (as shown in Tab. VI). Given the growing popularity of deep reasoning models such as OpenAI’s o1 and DeepSeek-R1, it is worth exploring whether these reasoning-enhanced models can achieve better forgetting generalisation on one-hop questions related to the forget data. To investigate this, we conduct experiments on DeepSeek-R1-Distill-Llama-8B using the ZSRE One-hop dataset. The model is first fine-tuned using LoRA with rank = 8, after which we apply various unlearning methods, updating only the LoRA parameters. As shown in Tab. XI, we have the following key observations:
[…]
In conclusion, deep reasoning models demonstrate improved one-hop unlearning performance. However, a substantial gap remains between the unlearning of forget data and one-hop data. Moreover, larger models face challenges in preserving utility while achieving effective unlearning, highlighting the need for further investigation in this area.
TABLE XIII: The impact of the percentage of perturbed tokens $K$ on Phi-1.3B. To balance forgetting and retention, we set $K=0.4$ in our main experiments.
TOFU Forget01 Dataset
Split Forget Rephrased Forget Retain MU↑ FRT↑
Metric RG↓ Pr↓ TR↑ RG↓ Pr↓ TR↑ RG↑ Pr↑ TR↑
K=0.1 47.19 37.96 58.74 40.94 32.32 59.51 82.04 88.47 48.45 51.71 1.31
K=0.2 44.98 35.66 60.50 40.20 29.35 59.83 79.18 87.27 47.96 51.55 1.38
K=0.3 43.93 32.75 61.64 39.75 28.41 61.38 80.70 87.34 47.41 51.92 1.43
K=0.4 43.38 30.58 63.76 39.70 27.69 61.71 79.17 87.21 47.98 51.21 1.44
K=0.5 42.99 30.29 64.87 38.41 25.09 63.41 78.42 87.42 48.08 51.42 1.50
K=0.6 42.85 30.02 65.96 37.63 26.71 64.05 77.77 85.76 46.67 51.55 1.50
K=0.7 42.49 30.51 66.06 36.76 26.04 65.54 77.24 85.29 46.36 52.17 1.54
K=0.8 42.19 28.39 67.52 37.00 25.21 66.56 77.73 85.36 46.31 51.82 1.56
K=0.9 41.89 29.31 67.21 36.98 26.13 66.85 77.17 85.08 46.19 50.39 1.50
K=1.0 41.57 27.84 68.46 36.74 24.5 68.08 76.28 84.63 46.24 50.24 1.54
7.3 Evaluation on More Diverse Datasets
To further evaluate the unlearning effectiveness, we conduct experiments on more challenging and diverse datasets—WMDP [23] and MUSE [24]—where knowledge is often distributed across long-form documents. For WMDP, we select 50 samples from BIO split as the forget set and 350 samples as the retain set. The target model is Vicuna-13B, fine-tuned using LoRA with rank = 8. For MUSE, we conduct experiments on BOOK split using MUSE-books-7B as the target model. To improve testing efficiency, we randomly sample 40 instances from the verbmem-forget split for unlearning and use the remaining 60 instances for retention. The average input length is 990.3 tokens, with a maximum length of 1031 tokens.
As shown in Tab. XII, RMU achieves the highest utility but at the cost of effective unlearning. This may be attributed to RMU’s design, which focuses on unlearning with limited samples rather than leveraging the entire forget set—thereby limiting its overall unlearning performance. In contrast, PerMU achieves a stronger forgetting effect, with improvements of 1.23% on WMDP and 3.24% on MUSE, and consistently attains the best FRT ratio. These results highlight the effectiveness of PerMU on larger models and longer-context scenarios.
TABLE XIV: Computational overhead of existing unlearning methods.
Metric Train step time (seconds) Memory Usage (MiB)
GA+KL 22.9342 35056
DPO+KL 47.1367 48132
NPO+KL 31.5571 36182
TV 14.2505 20442
ULD 32.4175 31014
WHP 24.1673 20442
PerMU 155.7616 52384
$\textsc{PerMU}^{\dagger}$ 24.3362 39054
7.4 Computational Overhead
We evaluated the computational overhead of existing methods from two perspectives: training step time and memory usage. The experiments were conducted on the Phi-1.3B model using the TOFU Forget01 dataset, with a batch size of 4 for 5 epochs. We report the total training step time and the maximum memory usage. As shown in Tab. XIV, although PerMU requires more training time, most of the overhead comes from computing the MSM to identify the top-$K$sensitive tokens. Notably, this step can be performed offline prior to training, significantly reducing both time and memory consumption during actual training. Moreover, our proposed fast variant, $\textsc{PerMU}^{\dagger}$, introduces no additional cost in training step time or memory usage, confirming its computational efficiency.
7.5 Percentage of Perturbed Tokens $K$
In PerMU, we identify the top-$K$ most sensitive tokens and apply perturbations to them. The choice of $K$ directly influences the unlearning effect. We analyze the impact of varying $K$ from 0.1 to 1.0 in increments of 0.1, where each value represents $K$of the tokens to be perturbed, while keeping all other hyper-parameters fixed. As shown in Tab. XIII, increasing $K$ consistently enhances the unlearning ability, with Forget Rouge improved by up to 5.62% (47.19 $\rightarrow$ 41.57) on the TOFU dataset. However, this improvement comes at the cost of reduced model utility. This trade-off is expected, as perturbing more tokens makes it harder for the model to recover and correctly predict the unlearning samples. Although the Forget Retain Trade-off (FRT) score peaks when $K=0.8$ to $0.9$, the performance on retained knowledge drops significantly—up to 5.76% in Rouge (82.04 $\rightarrow$ 76.28) and 3.84% in Probability (88.47 $\rightarrow$ 84.63). Such degradation may negatively impact the model’s utility on preserved knowledge. To strike a balance between forgetting and retention, we set $K=0.4$ in our main experiments, which achieves effective unlearning while preserving most of the model’s capacity to retain relevant knowledge.
TABLE XV: The impact of the perturbation ratio $P$. Experiments are conducted on the TOFU Forget01 dataset using Phi-1.3B. Different values of $P$ lead to varying trade-offs, we select $P=0.4$ as the optimal perturbation ratio.
Dataset Forget Rephrased Forget MU↑ FRT↑
Metric RG↓ Pr↓ TR↑ RG↓ Pr↓ TR↑
P=0.1 52.12 43.00 56.18 42.28 31.00 53.79 50.90 1.21
P=0.2 50.08 37.28 57.03 40.74 27.06 57.53 50.55 1.30
P=0.3 48.02 34.68 61.40 42.16 27.50 59.13 51.12 1.34
P=0.4 46.26 27.11 64.65 39.95 22.76 63.44 50.14 1.47
P=0.5 47.17 30.64 62.70 41.09 24.62 62.60 48.00 1.34
P=0.6 44.47 25.99 62.43 41.33 21.39 62.61 48.61 1.46
P=0.7 41.69 26.95 65.19 41.10 22.84 64.38 48.79 1.47
P=0.8 42.37 24.87 65.03 39.97 21.18 64.96 47.51 1.48
P=0.9 41.11 22.49 66.79 38.76 19.08 64.88 48.08 1.58
P=1.0 41.53 21.10 68.66 37.11 18.31 66.21 48.61 1.65
7.6 Perturbation Ratio $P$
We analyze the impact of the Perturbation Ratio $P$, which ranges from $0.1$ to $1.0$ in increments of $0.1$. The experiments are conducted on the TOFU Forget01 dataset using Phi-1.3B, keeping other parameters constant.
As shown in Tab. XV, even a small amount of noise ($P=0.1$) is sufficient to achieve a notable unlearning effect. Moreover, increasing the noise ratio further enhances the effectiveness of unlearning. Specifically, when $P=1.0$, almost all metrics on the Forget and Rephrased Forget data achieve their best values. This is expected, as higher noise levels make it harder for the model to recall related facts, resulting in a more fact-unrelated probability distribution and better unlearning performance.
TABLE XVI: The impact of the Tuning Coefficient $C$. Experiments are conducted on the TOFU Forget01 dataset using Phi-1.3B. Increasing $C$ results in an improved unlearning effect, but at the cost of decreased model utility. To balance model utility with effective unlearning, we select $C=0.1$ for our experiments.
Dataset Forget Rephrased Forget MU↑ FRT↑
Metric RG↓ Pr↓ TR↑ RG↓ Pr↓ TR↑
C=0.0 49.69 36.45 62.02 42.08 29.45 59.81 50.50 1.28
C=0.1 46.26 27.11 64.65 39.95 22.76 63.44 50.14 1.47
C=0.2 37.62 13.66 70.43 38.76 12.18 70.10 47.73 1.87
C=0.3 24.64 2.84 71.73 25.77 2.70 72.37 39.66 2.84
C=0.4 12.76 0.83 73.13 15.39 0.88 72.89 31.85 4.27
C=0.5 18.11 1.38 75.68 17.16 1.36 75.04 38.81 4.08
C=0.6 19.54 1.54 77.77 21.73 1.52 77.60 37.56 3.39
C=0.7 18.51 1.91 78.28 20.64 1.69 77.36 37.59 3.52
C=0.8 16.15 1.69 78.65 19.31 1.43 77.10 37.35 3.87
C=0.9 15.46 1.52 77.68 17.25 1.28 76.50 36.92 4.16
C=1.0 17.64 1.48 77.19 18.66 1.24 76.21 36.66 3.76
However, the model utility decreases by up to 3.61% when $P\geq 0.5$. This suggests that excessive noise can hinder the model’s sentence comprehension and increase uncertainty, unintentionally affecting irrelevant knowledge generation. The model utility is the highest when $P=0.3$. Since different values of $P$ result in varying trade-offs, we select $P=0.4$ as the optimal perturbation ratio. This choice is based on the fact that the FRT ratio at $P=0.4$ is higher than at $P=0.3$, while maintaining considerable model utility, indicating a better balance between unlearning performance and model utility. Therefore, we intuitively set $P=0.4$ for all experiments.
7.7 Tuning Coefficient $C$
We investigate the impact of the Tuning Coefficient $C$, which varies from $0.0$ to $1.0$ in increments of $0.1$. The experiments are conducted on the TOFU Forget01 dataset using Phi-1.3B, with all other parameters held constant. As shown in Tab. XVI, the best model utility is achieved when $C=0.0$, where only the corrupted-run probability distribution is used. While this setting maintains high model utility, the unlearning effect is insufficient, as the top-ranked token in the clean-run probability distribution is not fully suppressed. As $C$ increases, the unlearning effect improves, reaching its peak when $C=0.4$. Then it begins to fluctuate as $C$ continues to increase. Correspondingly, model utility decreases with larger values of $C$, which is expected, as higher $C$ values subtract more information from $p(y|x^{\prime})$, potentially disrupting the distribution of irrelevant knowledge. To balance model utility with effective unlearning, we select $C=0.1$ for our experiments.
7.8 Discrete-Token Level Perturbation
Apart from adding random noise to the subject token embeddings, PerMU can also be implemented by perturbing the subject words at the discrete-token level, denoted as $\textsc{PerMU}^{\dagger}_{dis}$. In this experiment, we evaluate the unlearning performance of $\textsc{PerMU}^{\dagger}_{dis}$ on the TOFU dataset while keeping all other parameters constant. The perturbation type is randomly chosen from deleting, altering, or adding letters to the subject words.
As shown in Tab. XVII, $\textsc{PerMU}^{\dagger}_{dis}$ exhibits exceptional unlearning capability, outperforming $\textsc{PerMU}^{\dagger}$ in Forget ROUGE across all datasets and models. Moreover, $\textsc{PerMU}^{\dagger}_{dis}$ achieves an absolute 19.26% Forget ROUGE and 4.64% Forget Probability on Forget05 when using LLaMA2-7B, surpassing $\textsc{PerMU}^{\dagger}$ by up to 11.16% (30.42$\rightarrow$19.26) and 11.65% (16.29$\rightarrow$4.64), respectively, though with a slightly lower model utility of 0.92%. However, when tested on Phi-1.3B, $\textsc{PerMU}^{\dagger}_{dis}$ does not consistently exhibit superior unlearning performance, and the model utility drops by up to 4.44% (50.94$\rightarrow$46.50) as the number of unlearning samples increases. Despite this, the FRT ratio of $\textsc{PerMU}^{\dagger}_{dis}$ still outperforms other baselines. In summary, perturbing subject words at the discrete-token level can also prevent the model from recalling the fact and generate fact-unrelated probability distributions, thus achieving unlearning. Both embedding-layer and discrete-token-level noise methods can achieve effective unlearning but result in different trade-offs. Given that the knowledge retention ability of $\textsc{PerMU}^{\dagger}_{dis}$ may decline as the amount of forgotten data increases, we choose to add noise at the embedding layer as a more promising alternative.
TABLE XVII: Experimental results for implementing perturbation at the discrete-token level to subject words, denoted as $\textsc{PerMU}_{dis}$. Surprisingly, $\textsc{PerMU}^{\dagger}_{dis}$ demonstrates a stronger unlearning effect on some datasets, particularly achieving 19.26% in Forget Rouge and 4.64% in Probability on the Forget01 dataset using LLaMA2-7B. These results highlight the adaptability of our approach across different dimensions.
Model TOFU Dataset Forget data Retain data Real Authors Real World MU↑ FRT↑
Metric RG↓ Pr↓ TR↑ RG↑ Pr↑ TR↑ RG↑ Pr↑ TR↑ RG↑ Pr↑ TR↑
Phi (1.3B) Forget01 $\textsc{PerMU}^{\dagger}$ 46.26 27.11 64.65 67.74 79.50 44.96 41.90 37.85 44.04 76.14 41.75 50.39 50.14 1.37
$\textsc{PerMU}^{\dagger}_{dis}$ 35.46 27.53 53.67 84.31 89.36 48.18 46.23 38.26 46.71 76.10 40.91 49.67 52.72 1.67
Forget05 PerMU 42.67 25.17 62.96 67.97 77.49 44.79 43.23 37.94 45.60 76.13 43.10 52.65 50.94 1.50
$\textsc{PerMU}^{\dagger}_{dis}$ 36.94 36.76 60.86 42.37 58.66 41.05 45.82 37.43 44.93 78.40 40.73 48.92 46.50 1.26
Forget10 $\textsc{PerMU}^{\dagger}$ 46.55 41.93 57.51 81.59 86.17 47.10 35.23 37.64 45.17 75.28 41.25 49.92 50.07 1.13
$\textsc{PerMU}^{\dagger}_{dis}$ 44.90 57.06 55.92 46.25 63.60 43.77 49.35 36.89 44.34 73.65 39.78 47.41 47.39 0.93
LLaMA (7B) Forget01 $\textsc{PerMU}^{\dagger}$ 30.42 16.29 74.11 86.67 88.55 43.51 92.80 51.82 66.34 89.17 49.28 63.02 65.06 2.79
$\textsc{PerMU}^{\dagger}_{dis}$ 19.26 4.64 73.92 83.72 89.11 42.39 91.00 51.19 66.05 87.46 48.78 62.66 64.14 5.37
Forget05 $\textsc{PerMU}^{\dagger}$ 33.66 39.23 64.39 83.63 88.24 41.89 91.30 52.57 68.21 89.60 49.77 63.94 64.89 1.78
$\textsc{PerMU}^{\dagger}_{dis}$ 30.39 36.55 65.45 72.08 79.45 41.19 91.50 52.61 68.67 88.32 50.15 64.52 63.38 1.89
Forget10 $\textsc{PerMU}^{\dagger}$ 44.60 47.58 67.25 94.14 94.68 41.23 90.30 51.66 66.56 88.75 48.48 62.10 64.80 1.41
$\textsc{PerMU}^{\dagger}_{dis}$ 40.32 48.68 62.45 70.72 79.47 41.56 91.50 53.81 70.03 87.89 50.86 65.92 63.93 1.44
TABLE XVIII: A comparison of different retain loss settings. w/o GDR refers to using the vanilla forget loss, while w/ KLR indicates the combination with KLR. Overall, $\textsc{PerMU}^{\dagger}$ effectively preserves the model’s utility.
Dataset Forget Rephrased Forget MU↑ FRT↑
Metric RG↓ Pr↓ TR↑ RG↓ Pr↓ TR↑
$\textsc{PerMU}^{\dagger}$ 44.60 47.58 67.25 32.72 37.99 65.96 64.80 1.59
w/o GDR 44.88 47.52 66.44 34.23 39.67 65.50 60.25 1.45
w/ KLR 41.54 45.43 67.45 31.92 38.15 66.44 60.55 1.54
7.9 Different Retain Loss
We investigate the impact of different retain loss functions while keeping other parameters fixed. The experiments are conducted on the TOFU Forget10 dataset using LLaMA2-7B. $\textsc{PerMU}^{\dagger}$ typically employs the forget loss combined with GDR, with a retain weight of $RW=1$. Here, w/o GDR refers to using the vanilla forget loss without any retain loss, while w/ KLR denotes the combination of the vanilla forget loss with KLR, applying the same retain weight. As shown in Tab. XVIII, using the vanilla forget loss achieves strong unlearning performance but may slightly impair model utility. While incorporating KLR can improve model utility, the enhancement is less significant compared to using GDR as the retain loss. Therefore, we primarily adopt GDR as the retain loss in our experiments.
8 Conclusions
In this paper, we advance LLM-based knowledge forgetting by shifting the focus from superficial forgetting to the more challenging task of implicit knowledge forgetting. To this end, we introduce UGBench, a comprehensive benchmark for evaluating implicit unlearning, and propose PerMU, a novel approach that addresses two core challenges identified by UGBench. Our results highlight the need for more generalised and thorough unlearning in LLMs. While automatic evaluation metrics offer reliable assessments of model performance, we plan to incorporate human evaluations in future work to further enhance the evaluation process. Overall, we believe our contributions offer meaningful insights into effective knowledge unlearning in NLP, with broader implications for related areas such as model editing, knowledge conflict resolution, reasoning, and other diverse fields.